For one deposit of 2000 EUR in ProCredit Bank you will have 30 EUR interest for 4 months. Find the interest rate for one year.
Answer by jorel1380(3719)

30=2000*4r/12=2000*r/3
2000r=90
r=90/2000=.045=4.5%
☺☺☺☺